In an election 10% of total voters didn't cast their votes and 20% votes were found invalid. A candidate got 40% of total votes and he won by 3600 votes. Find total number of voters in votin list?

Grade:8

3 Answers

Vijay Mukati
7 years ago
Dear Student,

From the given question, 30% of the total votes are not usefull for either of the candidate.
Out of remaining 70% of the valid votes, winning candidate got 40% of the votes and won by 3600 votes.
These means that the difference in percentage of votes to both the candidates is equal to 3600 votes
or 10% of total voters = 3600
or total voters = 3600*100/10 = 36000.

Thanks.
Seema
7 years ago
45000 is the answer.
The winning candidate got 40% of total and not 40% of valid votes.         
 Therefore 30% in all cannot be taken out of the calculation.  Please suggest.
Vijay Mukati
7 years ago
Sorry for the careless mistake.

Let the total no. of voters be x.
Out of these 0.9x will vote.
Now only 80% of these are valid. which means valid votes are 0.72x
Now wining candidate get 40% of total votes i.e. 0.4x and remaining valid vote i.e. 0.72x – 0.4x = 0.32x will be of opposite candidates.
Now according to given condition, 0.4x – 0.32x = 3600
or 0.08x = 3600 === x = 45000

Thanks.